OpenStudy (anonymous):

c). If f(x)={x*(sin|x|/x) for x is not equal to zero {0 for x=0 , then discuss continuity of f(x) at x=0.

OpenStudy (anonymous):

if the limit as x appraoches 0 and (f0) is the same then it is continuous: \[\lim_{x \rightarrow 0}=f(0)?\]

OpenStudy (anonymous):

plz write complete solution step by step

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{xsin|x|}{x}=f(0)?\] f(0)=0 from ur problem so: \[\lim_{x \rightarrow 0}\frac{xsin|x|}{x}=0?\] the x cancels out so: \[\lim_{x \rightarrow 0}\sin|x|=\sin(0)=0\] and 0=0 so then the limit does equal to point so it is continuous

OpenStudy (anonymous):

plz look at the question attached

OpenStudy (anonymous):

ok u have to use squeeze theorem

OpenStudy (anonymous):

u kno that sin of any angle has to b between -1 and 1 right? so: -1<sin(any angle)<1

OpenStudy (anonymous):

\[-1 \le \sin(anyangle) \le 1\]

OpenStudy (anonymous):

in this case the angle is: \[\frac{|x|}{x}\] so \[-1 \le \sin(\frac{|x|}{x}) \le 1\] now lets multiply each side by x so it becomes: \[-x \le x*\sin(\frac{|x|}{x}) \le x\] now the middle is the limit u want to solve. so lets solve the left and right side for the limit we want (0) \[\lim_{x \rightarrow 0} -x \le x*\sin(\frac{|x|}{x}) \le x\] becomes \[0 \le x*\sin(\frac{|x|}{x}) \le 0\] so it must be greater than or equal to 0 and less than or equal to 0 and the only value that can fulfil both of these is 0 so it must equal 0 so ur limit and point are equal and it is continuous

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}x*\sin(\frac{|x|}{x})=0\]

hero (hero):

Yo